OpenStudy (anonymous):

use the ratio test to determine if series is convergent for sigma 1 to infinity (n!)^2/(2n)!

OpenStudy (anonymous):

An = (n!)^2/(2n)! An+1 = ((n+1)!)^2/(2n+1)! and then abs (An+1/An)

OpenStudy (anonymous):

hi im still confused

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

i solved the problem and my final answer is n+1/2 and im not getting the right answer

OpenStudy (anonymous):

do you know how 2(n+1)! becomes( 2n+1)(2n+2)

OpenStudy (anonymous):

n! = n (n-1)(n-2)...1 (2n!) = 2(n!) = 2(n(n-1)(n-2)...1) (n+1)!= n+1((n+1)-1)(n+1)-2)...1) like that